OpenStudy (anonymous):

how would you identify the x intercepts of f(x)=25-x^2? No graphing calculator allowed.

OpenStudy (anonymous):

set 25-x^2=0

OpenStudy (anonymous):

and then factor (x+5)(x-5)?

OpenStudy (anonymous):

Take y=0 for finding x-intercepts. 0=25-x^2 Take x^2 to the other side, x^2=25 Or x=5

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so \[x=\pm 5\]

OpenStudy (anonymous):

\[x^2=25, x= \pm 5\]

OpenStudy (anonymous):

Yes that's right.

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

wait.. but it's a negative x..

OpenStudy (anonymous):

There are 2 solutions, positive and negative 5

OpenStudy (anonymous):

OH, you moved it to the other side, gotcha

OpenStudy (anonymous):

yeah no need to factor that one, pretty easy :)

OpenStudy (anonymous):

wasn't thinking straight, what a simple question that was.. hahah wow. -facepalm-

OpenStudy (anonymous):

It happens, because of our awesome education systems we tend to make easy things difficult sometimes.